OpenStudy (anonymous):

Calculate the molarity of 1.31 L of a solution containing 1.61 g of dissolved KBr.

OpenStudy (anonymous):

Molarity=Moles/Volume. Moles= Weight/Molecular weight= 1.61/119=0.0135 Molarity=0.0135/1.31=0.01